Find the dimensions of the rectangle that has a length 3 meters more than its width and a perimeter equal in value to its area?
kindly help me with these problem? thanks

Found 2 solutions by stanbon, Cromlix:
Answer by stanbon(75887)

Find the dimensions of the rectangle that has a length 3 meters more than its width and a perimeter equal in value to its area?
-----
width = w
length = w+3
-----
Perimeter = 2(width+length) = 2(w+w+3) = 4w+6
Area = width*length = w(w+3) = w^2+3w
-------
Equation:
Perimeter = Area
w^2 + 3w = 4w+6
w^2 - w -6 = 0
----
Factor:
(w-3)(w+2) = 0
------------------------
Positive Solution:
width = w = 3
length = w+3 = 6
=====================
Cheers,
Stan H.

Answer by Cromlix(4381)

Width = x
Length = x + 3
Perimeter = 2* width + 2*length
Area = width * length
Perimeter = Area
2*x + 2*(x + 3) = x*(x + 3)
2x + 2x + 6 = x^2 + 3x
x^2 + 3x - 4x - 6 = 0
x^2 - x - 6 = 0
(x + 2)(x - 3) = 0
x + 2 = 0
x = -2 (No answer x<0)
x - 3 = 0
x = 3.
So, width = 3m
length = 6m.
Proof:
Perimeter = Area
2*3 + 2*6 = 3*6
18m = 18m
Hope this helps.
:-)